OpenStudy (anonymous):

What would be the most logical first step to solve this quadratic equation? x2+2x+13=8

OpenStudy (anonymous):

what level is this? (that equation only has imaginary roots and, without stereotyping, most people who know about imaginary numbers can solve quadratics)

OpenStudy (anonymous):

just use ur quadtratic equatin and substitute.

OpenStudy (anonymous):

so would it be the square root of both sides .

OpenStudy (anonymous):

is it x^2?

OpenStudy (anonymous):

Did you make up this question yourself?

OpenStudy (anonymous):

no its algebra twilight math class .

OpenStudy (anonymous):

its not complicated, just put in the quadtratic formula, the values of Ax +Bx +C=0, where in ur case, its x^2 +2x +5 =0..... or am i missing something...

OpenStudy (anonymous):

Hmm. Well the first step would be to either set it to 0, or complete the square. It won't factorise, so I'd complete the square. But you'll still need to square root a negative number.

OpenStudy (anonymous):

i dont know how to do that sorry im not qood at math worse subject evr .

OpenStudy (anonymous):

Note \[\sqrt{-1} = i \]

OpenStudy (anonymous):

where do u see a negative value?

OpenStudy (anonymous):

nvm is see it now...

OpenStudy (anonymous):

{ -2 ± Sq Root ( 4 - 20 ) } / 2

OpenStudy (anonymous):

unless u use imaginary numbers, thats as simplified as it gets...

OpenStudy (amistre64):

subtract the right hand side from both sides of the equation would be the first step

OpenStudy (amistre64):

x2+2x+13 -8=8 -8 x^2 +2x +5 = 0

OpenStudy (amistre64):

the next step would be to "complete the square".... if you dont know the quadratic formula that is :) x^2 +2x +5 = 0 ; subtract 5 from both sides x^2 +2x + ___ = -5 + ____ we need to find a suitable number to fill in these blanks with; we need to create a "perfect" square trinomial that will factor into: (x+?)^2 we know that the product of a binomial and itself is: (a+b)(a+b) = a^2 +2ab + b^2 what we have is: a^2 +a(2__ ) + (___)^2

OpenStudy (anonymous):

I know the quadratic formula, but would still complete the square.

OpenStudy (amistre64):

if I recall correctly, completing the sqaure is how we get the quadratic formula :)

OpenStudy (anonymous):

Affirmative.

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

man im in university calc, and i dont even remember how to complete the square...havent done that since 9th grade haha

OpenStudy (amistre64):

(a+b)(a+b) = a^2 +2ab + b^2 what we have is: a^2 +a(2__ ) + (___)^2 we see that if we divide the midde term by 2a we will get a suitable "b" right? x^2 +2x +5 ....2x/2x = 1...b=1 b^2 = 1

OpenStudy (anonymous):

pellet, university calc. That's harcore.

OpenStudy (anonymous):

hardcore* :@

OpenStudy (amistre64):

x^2 +2x +1 = -5+1 (x+1)^2 = -4 already see a problem lol x+1 = +-sqrt(-4) x = -1 +- sqrt(-4)

OpenStudy (anonymous):

its also really hard, and i dont know why i need it if im in software engineering, but ok...

OpenStudy (anonymous):

Because Maths is everything.

OpenStudy (amistre64):

x = -1 +- 2i :) but thats a complex number and not a "real" number

OpenStudy (anonymous):

\[x^2 + 2x + 13 = 8 \] \[\iff (x+1)^2 = -4 \] \[\implies (x+1) = \pm \sqrt{-4} = \pm \sqrt{-1}\sqrt{4} = \pm 2i \] \[\implies x = -1 \pm 2i \]

OpenStudy (anonymous):

sorry had that typed out but realised you were explaining it so waited - but didn;t want all that typing to go to waste.

OpenStudy (amistre64):

lol ..... you did good ;)

OpenStudy (amistre64):

does that make any sense miranda?

OpenStudy (anonymous):

... :(

OpenStudy (anonymous):

im sure thats a yes ;)